# Towards Robust Infrared Small Target Detection: A Feature-Enhanced and Sensitivity-Tunable Framework

### v2

@@ Abstract
Recently, single-frame infrared small target (SIRST) detection technology has attracted wide-spread attention. However, due to the intrinsic feature scarcity in infrared small targets, precise segmentation of small targets from complex backgrounds remains a significant challenge. Different from most existing deep learning-based methods that focus on improving network architectures, we propose a feature-enhanced and sensitivity-tunable (FEST) framework, which is compatible with existing SIRST detection networks and further enhances their detection performance. The FEST framework improves the model’s robustness from two aspects: feature enhancement and target confidence regulation. For feature enhancement, on the one hand, we adopt a multi-scale fusion strategy, which can effectively improve the model’s perception and adaptability to multi-scale features of multi-size targets. On the other hand, we construct an edge enhancement difficulty mining (EEDM) loss based on the analysis of the task characteristics, which helps guide the network to continuously focus on challenging target regions and edge features during training. For target confidence regulation, we design an adjustable sensitivity (AS) strategy for network post-processing. This strategy not only enhances the adaptability of the network in complex scenarios, but also significantly improves the detection rate of infrared small targets while maintaining segmentation accuracy. Extensive experimental results show that our FEST framework can significantly enhance the performance of existing SIRST detection networks. Notably, the multi-scale direction-aware network (MSDA-Net) equipped with the FEST framework won the first prize in the “PRCV 2024 wide-area infrared small target detection competition”.
@@ 1 Introduction
Infrared imaging technology, with its excellent low-light imaging capabilities, demonstrates unique advantages in complex environment perception, driving the widespread application of infrared detection technologies [1, 2, 3, 4]. Among them, SIRST detection, as a fundamental technology in infrared detection technology, plays an important role in many fields, including target recognition and tracking [5, 6], early warning systems [7, 8], traffic surveillance [9, 10], and maritime search and rescue [11, 12].
@@
The research on SIRST detection can be divided into non-deep learning-based methods and deep learning-based methods. Non-deep learning-based SIRST detection methods mainly use the characteristics of infrared small target images for detection [13, 14, 15, 16, 17, 18, 19, 20]. These methods often employ traditional image processing techniques, including filtering, sparse representation, low-rank decomposition, and background modeling, to enhance target features while reducing background interference. They are usually optimized for specific scenarios and can achieve good detection results under specific imaging conditions. However, owing to the complexity and variability of practical scenarios, these traditional methods often suffer from many false detections or even failure. With the rapid development of deep learning, a large number of researchers have focused on leveraging the powerful representation capabilities of deep neural networks to extract richer, more efficient, and more robust features, which have been widely applied across various research fields [21, 22, 23, 24]. In light of current technological trends and growing application demands, deep learning-based SIRST methods [25, 26, 27, 28, 29, 30, 31, 32, 33] have achieved remarkable progress. This type of method usually uses deep learning models such as convolutional neural network (CNN) [34] and visual transformers (ViT) [35] to automatically learn the spatial and texture features of the target from the data, thereby improving detection performance. However, most current methods mainly focus on optimizing the network structure while ignoring other aspects that may improve the detection performance. Therefore, it is of great significance to design an efficient and robust SIRST detection framework to enhance the overall performance of networks in complex environments.
@@
Different from most existing deep learning-based methods that focus on improving network architectures, we propose a feature-enhanced and sensitivity-tunable (FEST) framework. The framework can be flexibly compatible with existing SIRST networks, thereby improving its robustness in complex scenarios. Specifically, in the aspect of the feature enhancement, a multi-scale fusion strategy is employed to improve the model’s perception and adaptability to multi-scale features of multi-size targets. By integrating information across multiple spatial scales, a comprehensive representation of multi-scale features of a single target is achieved, and adaptability to the scale variation in multiple targets is enhanced. In addition, to guide the network to pay more attention to challenging target regions and edge features during training, we propose an edge enhancement difficulty mining (EEDM) loss. This loss helps the model learn valuable features more effectively and greatly avoids the risk of small targets being submerged by the background. In the aspect of the target confidence regulation, to further enhance the adaptability of our framework to different scenarios, we propose an adjustable sensitivity (AS) strategy for post-processing. This strategy introduces the concepts of strong targets and weak targets. It can significantly improve the detection rate of infrared small targets while ensuring a small change in the false alarm rate and segmentation accuracy, thus having important practical application value. Extensive experimental results show that the proposed FEST framework can significantly enhance the performance of existing SIRST detection networks. The contributions of this manuscript can be summarized as follows:
@@
    A feature-enhanced and sensitivity-tunable (FEST) framework is constructed, which can be flexibly compatible with the existing SIRST network and effectively enhances their detection performance.
@@
    For feature enhancement, a multi-scale fusion strategy is adopted to improve the model’s perception and adaptability to multi-scale features of multi-size targets. In addition, an edge enhancement difficulty mining loss is constructed, which can guide the network to pay more attention to edge features and difficult target regions.
@@ 2.1 Non-deep learning-based SIRST detection methods
Non-deep learning infrared small target detection methods can be roughly divided into three categories: background suppression-based methods, human visual system-based methods and image data structure-based methods. Among them, background suppression-based methods include spatial domain filtering (e.g., Max-Median filter [13] and two-dimensional minimum mean square error filter [14]) and transform domain filtering (e.g., dual-tree complex wavelet [15] and bidimensional empirical mode decomposition (BEMD) [16]). Spatial domain filtering methods are computationally efficient but suffer performance degradation in complex scenes due to the difficulty in predefining effective templates. In contrast, transform domain methods offer stronger background modeling capabilities but incur high computational costs, limiting their practical applicability. The human visual system-based methods simulate the human eye’s perception of local contrast to detect salient objects in images. Representative approaches include the local contrast measure (LCM) [17] and multi-scale patch-based contrast measure (MPCM) [18]. These methods can effectively suppress large-area bright background interference and improve target saliency. However, they still tend to produce many false detections when strong noise or interference is present in the image. Image data structure-based methods formulate the detection task as a low-rank and sparse matrix separation problem, leveraging the non-local self-similarity of the background and the sparsity of the target to achieve foreground-background separation. Representative approaches include the infrared patch-image model (IPI) [19] and the non-negative infrared patch-image model based on partial sum minimization of singular values (NIPPS) [20]. These methods generally offer effective separation performance but remain sensitive to strong edges in the background and are computationally expensive.
Figure 1: Overall structure of the FEST framework. SIRST Network denotes the selected existing SIRST networks. Size $O_{1}$, …, Size $O_{n-1}$, and Size $O_{n}$ denote the actual sizes of the original training images. Size 1, Size 2, and Size 3 denote the input resolutions corresponding to each model. Size $O_{t1}$, …, Size $O_{t(m-1)}$, and Size $O_{tm}$ denote the actual sizes of the original test images.
@@ 2.2 Deep learning-based SIRST detection methods
Different from non-deep learning-based methods, deep learning-based methods learn the features of infrared small targets in a data-driven manner [36, 37]. Specifically, Dai et al. propose the asymmetric context module (ACM) [25] and the attention local contrast network (ALCNet) [26], which for the first time introduced the local contrast idea in traditional methods into the deep learning-based method. To further explore the relationship between local areas, Yu et al. successively proposed the multi-scale local contrastive learning network (MLCL-Net) [27] and the improved local contrastive learning network (ALCL-Net) [28], which enhanced the network’s perception ability of infrared small targets. Subsequently, Li et al. proposed a densely nested attention network (DNANet) [29] based on the UNet++ architecture , aiming to alleviate the problem of target information loss caused by downsampling operations in infrared small target detection. At the same time, Wu et al. proposed a U-Net nested network (UIU-Net) [30], which achieves richer multi-scale feature expression by nesting a lightweight U-Net network in the backbone U-Net. In recent years, Zhao et al. further emphasize the gradient information and directional characteristics of infrared small target images and propose GGL-Net [31] and MSDA-Net [32] respectively. GGL-Net introduces amplitude images at multiple stages of network training, prompting the network to pay more attention to discriminative features. MSDA-Net further promotes the refinement of infrared small target segmentation by emphasizing the attention to multi-directional high-frequency features. In addition, Zhang et al. proposed an infrared shape network (ISNet) [33] to improve the accuracy of target recognition by emphasizing the extraction of shape features in infrared images. However, most current methods focus on optimizing the network structure while ignoring other aspects that may improve the detection performance. Therefore, we aim to build an efficient and robust SIRST detection framework to further improve the performance of existing SIRST detection networks.
@@ 3.1 Overall Framework
From fig. 1, the proposed scheme consists of two parts: the model training phase and the model inference phase. In the model training phase, first, we perform multi-scale transformations on the original training images, adjusting them to multiple resolutions through scaling operations. Then, the images at different scales are individually fed into the corresponding basic models for training. The basic models can be any existing SIRST network. Finally, to further improve the detection accuracy of the network, we use the proposed EEDM Loss as the loss function to optimize the network. In the model inference phase, first, we perform multi-scale transformations on the original training images and input the transformed images into the corresponding multi-scale models for inference. Next, the feature maps generated by multiple scale models are uniformly resampled to the original test image size to ensure feature alignment. Then, the multi-scale fusion strategy is used to integrate the prediction results of multiple scale models. This strategy can improve the model’s perception of different target scales and enhance its robustness in complex environments. Finally, the proposed AS strategy is used to post-process the fused results. This strategy can significantly improve the detection rate of infrared small targets while ensuring the false alarm rate and segmentation accuracy.
@@ 3.2 Multi-scale fusion strategy
In the SIRST detection task, the target scale is affected by factors such as imaging scene, shooting distance and sensor equipment, and can usually be divided into point targets, spot targets and extended targets [4, 32]. In complex scenes, a single-scale network model struggles to handle small targets of various sizes simultaneously, often resulting in missed or false detections. Additionally, considering that this study aims to build a detection framework compatible with the existing SIRST network, we seek to enhance detection performance without altering the network structure. Therefore, we adopt a multi-scale fusion strategy to enhance the network’s adaptability to different scales targets by integrating feature information from multiple scale models, effectively alleviating the shortcomings of a single-scale model.
@@
The multi-scale model fusion strategy is conducive to fully mining the semantic features and detail features of multi-size targets at different scales, making the model more accurate and reliable in identifying and locating targets. In addition, practical detection scenarios are often more complex and changeable, so fusing models of different scales can help improve the mode’s robustness in complex backgrounds. In practical applications, the fusion scales and the number of models can be reasonably chosen according to specific task requirements to achieve an optimal balance between detection performance and inference efficiency.

### v3

@@ Abstract
Recently, single-frame infrared small target (SIRST) detection technology has attracted widespread attention. Different from most existing deep learning-based methods that focus on improving network architectures, we propose a feature-enhanced and sensitivity-tunable (FEST) framework, which is compatible with existing SIRST detection networks and further enhances their detection performance. The FEST framework improves the model’s robustness from two aspects: feature enhancement and target confidence regulation. For feature enhancement, we employ a multi-scale fusion strategy to improve the model’s perception to multi-scale features of multi-size targets, and design an edge enhancement difficulty mining (EEDM) loss to guide the network to continuously focus on challenging target regions and edge features during training. For target confidence regulation, an adjustable sensitivity (AS) strategy is proposed for network post-processing. This strategy enhances the model’s adaptability in complex scenarios and significantly improves the detection rate of infrared small targets while maintaining segmentation accuracy. Extensive experimental results show that our FEST framework can effectively enhance the performance of existing SIRST detection networks. The code is available at https://github.com/YuChuang1205/FEST-Framework.
@@ 1 Introduction
Infrared imaging technology, with its excellent low-light imaging capabilities, demonstrates unique advantages in terms of complex environment perception, driving the widespread application of infrared detection technologies [han2022kcpnet, gao2024defense, terren2023detection, yu2024easy]. Among them, SIRST detection, as a fundamental technology, plays an important role in many fields, including target recognition and tracking [huang2023anti, ying2025infrared], early warning systems [xu2023multiscale, zhao2022single], traffic surveillance [ying2025visible, odat2017vehicle], and maritime search and rescue [zhao2024infrared, zhang2022rkformer].
@@
The research on SIRST detection can be divided into non-deep learning-based methods and deep learning-based methods. Non-deep learning-based SIRST detection methods mainly use the characteristics of infrared small target images for detection [arce1987theoretical, tomasi1998bilateral, li2010robust, chen2014novel, chen2013local, wei2016multiscale, gao2013infrared, dai2017non]. These methods often employ traditional image processing techniques, including filtering, sparse representation, low-rank decomposition, and background modeling, to enhance target features while reducing background interference. They are usually optimized for specific scenarios and can achieve good detection results under specific imaging conditions. However, owing to the complexity and variability of practical scenarios, these traditional methods often suffer from many false detections or even failures. With the rapid development of deep learning, many researchers have focused on leveraging the powerful representation capabilities of deep neural networks to extract richer, more efficient, and more robust features, which have been widely applied across various research fields [yu2023feature, tian2018detection, yu2022precise, zhang2023motor, raja2025bayesian]. In light of current technological trends and growing application demands, deep learning-based SIRST methods [dai2021asymmetric, dai2021attentional, yu2022infrared, yu2022pay, li2022dense, wu2022uiu, zhao2023gradient, zhao2024multi, zhang2022isnet] have achieved remarkable progress. This type of method usually uses deep learning models such as convolutional neural networks (CNNs) [lecun1998gradient] and Vision Transformers (ViT) [dosovitskiy2020image] to automatically learn the spatial and textural features of the target from the data, thereby improving detection performance. However, most current methods mainly focus on optimizing the network structure while ignoring other aspects that may improve the detection performance. Therefore, it is of great significance to design an efficient and robust SIRST detection framework to enhance the overall performance of networks in complex environments.
@@
Different from most existing deep learning-based methods that focus on improving network architectures, we propose a feature-enhanced and sensitivity-tunable (FEST) framework. The framework can be flexibly compatible with existing SIRST networks, thereby improving its robustness in complex scenarios. Specifically, in the aspect of the feature enhancement, a multi-scale fusion strategy is employed to improve the model’s perception to multi-scale features of multi-size targets. By integrating information across multiple spatial scales, the framework enables comprehensive multi-scale feature representation of a single target and enhances adaptability to scale variations among multiple targets. In addition, to guide the network to pay more attention to challenging target regions and edge features during training, we propose an edge enhancement difficulty mining (EEDM) loss. This loss helps the model learn valuable features more effectively and greatly avoids the risk of small targets being submerged by the background. In the aspect of the target confidence regulation, to further enhance the adaptability of our framework to different scenarios, we propose an adjustable sensitivity (AS) strategy for post-processing. This strategy introduces the concepts of strong targets and weak targets. It can significantly improve the detection rate of infrared small targets while ensuring a small change in the false alarm rate and segmentation accuracy, thus having important practical application value. Extensive experimental results show that the proposed FEST framework can significantly enhance the performance of existing SIRST detection networks. Notably, the multi-scale direction-aware network (MSDA-Net) equipped with the FEST framework won the first prize in the “PRCV 2024 wide-area infrared small target detection competition”. The contributions of this manuscript can be summarized as follows:
@@
    A feature-enhanced and sensitivity-tunable (FEST) framework is constructed, which can be flexibly compatible with the existing SIRST networks and effectively enhances their detection performance.
@@
    For feature enhancement, a multi-scale fusion strategy is adopted to improve the model’s perception to multi-scale features of multi-size targets. In addition, an edge enhancement difficulty mining loss is constructed, which can guide the network to pay more attention to edge features and challenging target regions.
@@ 2.1 Non-deep learning-based SIRST detection methods
Non-deep learning infrared small target detection methods can be roughly divided into three categories: background suppression-based methods, human visual system-based methods and image data structure-based methods. Among them, background suppression-based methods include spatial domain filtering (e.g., the Max-Median filter [arce1987theoretical] and two-dimensional minimum mean square error filter [tomasi1998bilateral]) and transform domain filtering (e.g., the dual-tree complex wavelet [li2010robust] and bidimensional empirical mode decomposition (BEMD) [chen2014novel]). Spatial domain filtering methods are computationally efficient but suffer performance degradation in complex scenes due to the difficulty in predefining effective templates. In contrast, transform domain methods offer stronger background modeling capabilities but incur high computational costs, limiting their practical applicability. Human visual system-based methods simulate the human eye’s perception of local contrast to detect salient objects in images. Representative methods include the local contrast measure (LCM) [chen2013local] and multi-scale patch-based contrast measure (MPCM) [wei2016multiscale]. These methods can effectively suppress large-area bright background interference and improve target saliency. However, they still tend to produce many false detections when strong noise or interference is present in the image. Image data structure-based methods formulate the detection task as a low-rank and sparse matrix separation problem, leveraging the non-local self-similarity of the background and the sparsity of the target to achieve foreground-background separation. Representative methods include the infrared patch-image model (IPI) [gao2013infrared] and the non-negative infrared patch-image model based on partial sum minimization of singular values (NIPPS) [dai2017non]. These methods generally offer effective separation performance but remain sensitive to strong edges in the background and are computationally expensive.
Figure 1: Overall structure of the FEST framework. “SIRST Network” denotes the selected existing SIRST networks. Size $O_{1}$, …, Size $O_{n-1}$, and Size $O_{n}$ denote the actual sizes of the original training images. Size 1, Size 2, and Size 3 denote the input resolutions corresponding to each model. Size $O_{t1}$, …, Size $O_{t(m-1)}$, and Size $O_{tm}$ denote the actual sizes of the original test images.
@@ 2.2 Deep learning-based SIRST detection methods
Different from non-deep learning-based methods, deep learning-based methods learn the features of infrared small targets in a data-driven manner [zhao2024refined, yu2024lr]. Specifically, Dai et al. propose the asymmetric context module (ACM) [dai2021asymmetric] and the attention local contrast network (ALCNet) [dai2021attentional], which for the first time introduced the local contrast idea in traditional methods into the deep learning-based method. To further explore the relationship between local areas, Yu et al. successively proposed the multi-scale local contrastive learning network (MLCL-Net) [yu2022infrared] and the improved local contrastive learning network (ALCL-Net) [yu2022pay], which enhanced the network’s perception ability of infrared small targets. Subsequently, Li et al. proposed a densely nested attention network (DNANet) [li2022dense] based on the UNet++ architecture , aiming to alleviate the problem of target information loss caused by downsampling operations in infrared small target detection. At the same time, Wu et al. proposed a U-Net nested network (UIU-Net) [wu2022uiu], which achieves richer multi-scale feature expression by nesting a lightweight U-Net network in the backbone U-Net. In recent years, Zhao et al. further emphasize the gradient information and directional characteristics of infrared small target images and propose GGL-Net [zhao2023gradient] and MSDA-Net [zhao2024multi] respectively. GGL-Net introduces amplitude images at multiple stages of network training, prompting the network to pay more attention to discriminative features. MSDA-Net further promotes the refinement of infrared small target segmentation by emphasizing the attention to multi-directional high-frequency features. In addition, Zhang et al. proposed an infrared shape network (ISNet) [zhang2022isnet] to improve the accuracy of target recognition by emphasizing the extraction of shape features in infrared images. However, most current methods focus on optimizing the network structure while ignoring other aspects that may improve the detection performance. Therefore, we aim to build an efficient and robust SIRST detection framework to further improve the performance of existing SIRST detection networks.
@@ 3.1 Overall Framework
From Fig. 1, the proposed scheme consists of two parts: the model training phase and the model inference phase. In the model training phase, first, we perform multi-scale transformations on the original training images, adjusting them to multiple resolutions through scaling operations. Then, the images at different scales are individually fed into the corresponding basic models for training. The basic models can be any existing SIRST networks. Finally, to further improve the detection accuracy of the network, we use the proposed EEDM loss as the loss function to optimize the network. In the model inference phase, first, we perform multi-scale transformations on the original test images and input the transformed images into the corresponding multi-scale models for inference. Next, the feature maps generated by multiple scale models are uniformly resampled to the original test image size to ensure feature alignment. Then, the multi-scale fusion strategy is used to integrate the prediction results of multiple scale models. This strategy can improve the model’s perception of different target scales and enhance its robustness in complex environments. Finally, the proposed AS strategy is used to post-process the fused results. This strategy can significantly improve the detection rate of infrared small targets while ensuring the false alarm rate and segmentation accuracy. To further illustrate the implementation of the FEST framework, the training and inference phases are presented in Algorithm 1 and Algorithm 2.
Algorithm 1 FEST Framework (Training phase)
1:Resize set $\{r_{1},r_{2},r_{3}\}$; Batch infrared images $I_{i=1}^{{N_{b}}}$; Batch ground-truth masks $M_{i=1}^{{N_{b}}}$; Network $\varphi$ (including $\varphi_{r1}$}, $\varphi_{r2}$}, $\varphi_{r3}$} for different input image sizes); Learning rate $l_{r}$.
2:Updated parameters $\{{\Theta_{{\rm{r1}}}},{\Theta_{r2}},{\Theta_{r3}}\}$
3:Initialize: Give the initial value to $\varphi_{r1}$}, $\varphi_{r2}$}, $\varphi_{r3}$}, and $l_{r}$.
4:Step 1: For each batch of data $\{(I,M)\}_{i=1}^{{N_{b}}}$, resize it to $\{r_{1},r_{2},r_{3}\}$:
$\{({I_{r1}},{M_{r1}})\}_{i=1}^{{N_{b}}}\;\leftarrow\;{Re_{r1}}(\{(I,M)\}_{i=1}^{{N_{b}}})$
$\{({I_{r2}},{M_{r2}})\}_{i=1}^{{N_{b}}}\;\leftarrow\;{Re_{r2}}(\{(I,M)\}_{i=1}^{{N_{b}}})$
$\{({I_{r3}},{M_{r3}})\}_{i=1}^{{N_{b}}}\;\leftarrow\;{Re_{r3}}(\{(I,M)\}_{i=1}^{{N_{b}}})$
5:Step 2: For each batch of data with different input sizes $\{({I_{r}},{M_{r}})\}_{i=1}^{{N_{b}}}$, data augmentation is performed to obtain $\{({I^{\prime}_{r}},{M^{\prime}_{r}})\}_{i=1}^{{N_{b}}}$.
6:Step 3: Feed each batch of the infrared images $\{\{{I^{\prime}_{r1}}\}_{i=1}^{{N_{b}}},\{{I^{\prime}_{r2}}\}_{i=1}^{{N_{b}}},\{{I^{\prime}_{r3}}\}_{i=1}^{{N_{b}}}\}$, into the corresponding network to obtain prediction maps $\{\{{P_{r1}}\}_{i=1}^{{N_{b}}},\{{P_{r2}}\}_{i=1}^{{N_{b}}},\{{P_{r3}}\}_{i=1}^{{N_{b}}}\}$:
$\{{P_{r1}}\}_{i=1}^{{N_{b}}}\;\leftarrow\;{\varphi_{r1}}(\{{I_{r1}}\}_{i=1}^{{N_{b}}})$
$\{{P_{r2}}\}_{i=1}^{{N_{b}}}\;\leftarrow\;{\varphi_{r2}}(\{{I_{r2}}\}_{i=1}^{{N_{b}}})$
$\{{P_{r3}}\}_{i=1}^{{N_{b}}}\;\leftarrow\;{\varphi_{r3}}(\{{I_{r3}}\}_{i=1}^{{N_{b}}})$
7:Step 4: Based on formulas (1)–(6), compute the EEDM loss using the prediction maps $\{\{{P_{r1}}\}_{i=1}^{{N_{b}}},\{{P_{r2}}\}_{i=1}^{{N_{b}}},\{{P_{r3}}\}_{i=1}^{{N_{b}}}\}$ and the ground-truth masks $\{\{{M^{\prime}_{r1}}\}_{i=1}^{{N_{b}}},\{{M^{\prime}_{r2}}\}_{i=1}^{{N_{b}}},\{{M^{\prime}_{r3}}\}_{i=1}^{{N_{b}}}\}$ to obtain $L_{r1}$, $L_{r2}$, $L_{r3}$.
8:Step 5: Based on the optimizer, update the network parameters $\{{\Theta_{{\rm{r1}}}},{\Theta_{r2}},{\Theta_{r3}}\}$:
${\Theta_{r1}}\;\leftarrow\;{\Theta_{r1}}\;-\;l_{r}{\nabla_{{\Theta_{r1}}}}{L_{r1}}$
${\Theta_{r2}}\;\leftarrow\;{\Theta_{r2}}\;-\;l_{r}{\nabla_{{\Theta_{r2}}}}{L_{r2}}$
${\Theta_{r3}}\;\leftarrow\;{\Theta_{r3}}\;-\;l_{r}{\nabla_{{\Theta_{r3}}}}{L_{r3}}$
9:Step 6: Repeat Step 1-5 until the training is completed.
10:Return: $\Theta_{{\rm{r1}}}$}, $\Theta_{r2}$}, $\Theta_{r3}$}.
@@ 3.2 Multi-scale fusion strategy
In the SIRST detection task, the target scale is affected by factors such as the imaging scene, shooting distance and sensor equipment, and can usually be divided into point targets, spot targets and extended targets [yu2024easy, zhao2024multi]. In complex scenes, a single-scale network model struggles to handle small targets of various sizes simultaneously, often resulting in missed or false detections. Additionally, considering that this study aims to build a detection framework compatible with the existing SIRST networks, we seek to enhance the detection performance without altering the network structure. Therefore, we adopt a multi-scale fusion strategy to enhance the network’s adaptability to different scales targets by integrating feature information from multiple scale models, effectively alleviating the shortcomings of a single-scale model.
@@
The multi-scale model fusion strategy is conducive to fully mining the semantic features and detail features of multi-size targets at different scales, making the model more accurate and reliable in identifying and locating targets. In addition, practical detection scenarios are often more complex and changeable, so fusing models of different scales can help improve the model’s robustness in complex backgrounds. In practical applications, the fusion scales and the number of models can be reasonably chosen according to specific task requirements to achieve an optimal balance between detection performance and inference efficiency.